# Stability analysis of power-law cosmological models

### v1

@@ Abstract
In this paper, we revisit the stability analysis of power-law models, focusing on an alternative approach that differs significantly from the standard approaches used in studying power-law models. In the standard approach, stability is studied by reducing the system of background FRW equations to a one-dimensional system for a new background variable $X$ in terms of the number of e-foldings. However, we rewrote the equations, including $H$ also into the system and went on to do the calculations up to second order. We demonstrate by computing the deviations from the power-law exact solution to second-order and show that power-law contraction is never an attractor, regardless of parameter values. Our analysis shows that while first-order corrections align with existing interpretations, second-order corrections introduce significant deviations that cannot be explained by a simple time shift that explains the first-order diverging terms. We also support our claim with numerical results. This new insight has broader implications for the study of attractor behaviour of differential equation solutions and raises questions about the stability of scenarios like the ekpyrotic bounce driven by an exponential potential.
@@ I Introduction
Cosmic inflation and the bounce paradigm represent two pivotal concepts in modern cosmology, offering distinct perspectives on the early universe’s evolution. Proposed by Alan Guth in 1981 [1], cosmic inflation postulates a rapid expansion phase that addresses the horizon problem and the flatness problem. On the other hand, the bounce paradigm challenges the classical Big Bang singularity, proposing a scenario where the universe transitions from a contracting phase to an expanding one without encountering a singularity [2, 3, 4, 5]. This concept has profound implications for cosmological models, providing alternative explanations for the universe’s origins. Bounce cosmology, though less explored compared to inflation, has garnered significant interest among researchers seeking to understand the early universe’s dynamics. Current studies in cosmology emphasize the refinement of both inflationary and bounce models, along with the search for observational evidence to support or refute these paradigms [6, 7]. For a comprehensive understanding of cosmological theories and observations, refer to relevant literature and books [8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
One of the early models of inflation is the elegant power-law inflation proposed by F. Lucchi and S. Matarrese [17]. The model assumes the evolution of the scale factor to be power-law ($a\propto t^{p}$, where $p$ is a positive constant greater than $1$). This is one of the simplest models of inflation that withstood even the latest observational tests [12, 18]. Power-law inflation and its spin-offs [19, 20, 21, 22, 23] are widely studied even today and form the basis of many cosmological models of early Universes which include models of inflation and cosmic bounce, and of late time acceleration. One of the most essential requirements of early Universe models is their ability to solve fine-tuning problems. Hence, a successful model must have the evolution of the Universe to be an attractor. For this reason, the stability analysis of both background evolution and the evolution of perturbations is important. In this paper, we reexamine the stability of the background solution of the power-law model. Indeed, it is a topic that has been studied widely in depth and breadth [24, 19, 25, 26, 27, 28, 29, 2]. However, the earlier studies, where they converted the FRW equations in time to equations in number of e-foldings for the two quantities defined as $X={\kappa\dot{\phi}}/{(\sqrt{6}H)}$ and $Y=\kappa\sqrt{V}/(\sqrt{3}H)$ [27] missed the point that, in doing so, we might not get the evolution of relevant quantities such as scalar field, Hubble parameter etc., in time. The evolution of these quantities in number of e-foldings is not enough to talk about the stability of their evolution. Instead one must look into the quantities of interest.
@@
One can note that it is impossible to write these background quantities as a function of the terms $X$ and $Y$. Our analysis, where we consider equations in time, redefines the conditions the parameters must satisfy for a stable expansion and contraction. Previously, there were studies in the literature in this regard, that is, studying the perturbations in time [29, 2]. However, they did the analysis to first-order and came to the conclusion that the perturbation would only result in a shift in time. From our numerical results, we conclude that this is not the case. Further, we analyzed the perturbations to the second-order and confirmed our conclusion.
@@
For expansion, that is for an inflationary evolution; though the analysis and results differ, the conclusion regarding the stability of the power-law expansion being an attractor remains the same without any change. However, the conditions for a stable contraction differ drastically. That is, our work is significant for the ekpyrotic bouncing model driven by the exponential potential.
@@
The outline of this paper is as follows: in the next section, we review the power-law model of inflation and the conditions for its stability. Then, we analytically study the background dynamics, giving importance to the stability of background evolution. Where we also compare our numerical results with analytical results. Subsequently, we show that $X$ is an attractor in N. Finally, in conclusion, we review the results and discuss their importance in the current studies of inflation, bouncing models and stability analysis in general.
@@ II Power-law models
where the constant $a_{0}$ is positive, the origin of $t$ has been chosen to coincide with $a(t)=0$  [^2] Since we do not expect general relativity to hold close to $a(t)=0$, we situate our discussion far enough from that point and the constant $p$ may in general be any real number. Neglecting the uninteresting constant scale factor of $p=0$, consider first the case $p>0$. Then, this scale factor will represent an expansion for $t>0$ and a contraction for $t<0$. For $p<0$, it will be the other way around. In this paper, we shall only consider cases with $p>0$.
@@
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa^{2}}R-\frac{1}{2}g^{ab}\nabla_{a}\phi\nabla_{b}\phi-V_{0}e^{\lambda\kappa\left(\phi-\phi_{0}\right)}\right]$ (3)
where general relativity is minimally coupled to a scalar field with an exponential potential with the constants $V_{0}$, $\lambda$ and $\phi_{0}$ as free parameters. The exponential potential allows a smooth inflationary expansion and perturbations to have a scale-invariant spectrum. where $t$ is the cosmic time and $a(t)$ is the scale factor. The Friedmann equations are given by
@@
$a(t)=a_{0}(C\,+\,q_{p}\,t)^{p};\quad\phi(t)=-\frac{\sqrt{2\,p}\,\ln\,\left(\mathit{q_{p}}t+C\right)}{\kappa}+\phi_{0}$ (7)
@@
There exists a relation between $p$ and $\lambda$, $p=2/\lambda^{2}$. We can obtain a contracting solution in two ways: either by assuming $q_{p}$ to be $-1$ and $C$ to take a large positive value, and also here, $t>0$. Now, this can also be achieved by assuming $t$ to take negative values ($\infty$- ∞ to 00), here again, $q_{p}=-1$. In the remaining part of the paper, we take $q_{p}=-1$, $C=0$, and $t$ takes negative values for a contracting solution. For an expanding solution $q_{p}=1$, $C=0$ and $t>0$ are assumed.
@@
$\frac{dX}{dN}$ $=-3X\,Y^{2}-\lambda\sqrt{\frac{3}{2}}Y^{2}$ (8)
11 11 $=X^{2}+Y^{2}$ (9)
@@
Now, we can see that $X=X_{p}=-\lambda/\sqrt{6}$ and $Y=Y_{p}=\pm\sqrt{1-\lambda^{2}/6}$ is a fixed point, i.e., for this point 8 and 9 are satisfied, and the derivatives of both $X$ and $Y$ are zero. If the evolution exactly falls at these points, it remains there indefinitely. Also, these are the very power-law solutions we obtained earlier, 7, where
@@
$V\left(\phi_{p}\right)$ $=V_{0}\,\left(\mathit{q_{p}}t+C\right)^{-2},\;\mbox{where}\;{V_{0}=M_{p}^{2}{q_{p}^{2}}\,p\,\left(3p-1\right)}.$ (14)
@@
We can clearly see that $Y_{p}$ is imaginary for $\lambda^{2}>6$. This is not unexpected as we need a negative value for $V$ in situations like bounce. Also, the value of $Y_{p}$ is not important; in the relevant equations and definitions, we don’t have $Y_{p}$; instead, what appears is $Y_{p}^{2}$.
Taking the Taylor expansion of 10 to the first order, we have
@@
So, if $\lambda^{2}<6$, the fixed point solutions are attractor solutions. This argument seems elegant. Also, it was verified numerically for inflationary solutions [24, 19]. For contracting solutions, the ekpyrotic case with $\lambda^{2}>6$ is considered an attractor [5]. For contracting solutions, we can argue that if expanding solutions are repellers, the corresponding contracting solutions would be attractors [36]. That is, $\lambda^{2}>6$ is the required condition for a power-law contraction to be an attractor. In the next section, we try to revisit the stability analysis.
III Re-examining the stability of power-law solutions
In the previous section, we obtained fixed-point power-law solutions and the conditions for these solutions to be attractors. However, the quantities that are studied are $X$ and $Y$. But the relevant quantities we have to study in order to see the nature of the evolution of the Universe in this model are $H$, $\phi$ and $\dot{\phi}$. Also, the analysis gives the conditions for the solutions $X$ and $Y$, defined in the previous section, to be stable when evolved in $N$. Note that those calculations couldn’t explicitly show the evolution of the Hubble parameter and the scalar field in time. Hence, we can argue that the analysis is not complete. Hence, we don’t think the stability of the evolution of solutions $X$ and $Y$ in $N$ guarantees that the evolution of $H$, $\phi$ and other relevant quantities will be stable in time. This possibility was explored in the classic paper [29]. In this section, we proceed to reproduce the solutions obtained by [29], analytically and then verify them numerically.
@@ III Re-examining the stability of power-law solutions
$\frac{dX}{dt}$ $=3HX^{3}+\frac{H\sqrt{6}\lambda X^{2}}{2}-\frac{H\sqrt{6}\lambda}{2}-3HX$
$=f_{1}(X,H)$ (20)
$\frac{dH}{dt}$ $=-3X^{2}H^{2}=f_{2}(X,H)$ (21)
The relevant solutions, as shown in 14 and 16 are $X=X_{p}$ and the Hubble parameter, which is not a constant in time or fixed point but a function of time, $H=H_{p}=p/t$ and $X=X_{p}=-\lambda/\sqrt{6}$. Now, Taylor expanding on the exact solution to the first order; we have
$\frac{d}{dt}\{X_{p}+\Delta X\}$ $=f_{1}(X_{p},H_{p})+\partial_{X}{f_{1}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta{X}+\partial_{H}{f_{1}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta H$ (22)
$\frac{d}{dt}\{H_{p}+\Delta H\}$ $=f_{2}(X_{p},H_{p})+\partial_{X}{f_{2}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta{X}+\partial_{H}{f_{2}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta H$ (23)
@@
$\dot{x}=\begin{pmatrix}\Delta{x_{1}}\\ \Delta{x_{2}}\end{pmatrix}^{\{.\}}\approx\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}\begin{pmatrix}\Delta x_{1}\\ \Delta x_{2}\end{pmatrix}$ (24)
@@
$x_{1}=\Delta X$, $x_{2}=\Delta H$ and
@@
$\Delta{X}$ $=$ ${D_{1}}t^{\frac{\lambda^{2}-6}{\lambda^{2}}}$ (26)
$\Delta{H}$ $=$ $\left[\frac{2\,D_{1}\sqrt{6}t^{\frac{\lambda^{2}-6}{\lambda^{2}}}}{t\lambda(\lambda^{2}-3)}+\frac{D_{2}}{(-p)}\frac{(-p)}{t^{2}}\right]$ (27)
@@
$\frac{\Delta N}{N_{p}}=1+D_{1}\frac{\lambda^{3}\sqrt{6}t^{\frac{\lambda^{2}-6}{\lambda^{2}}}}{(\lambda^{4}-9\lambda^{2}+18)\,\mbox{ln}(|t|)}-D_{2}\frac{\lambda^{2}}{2\,t\,\mbox{ln}(|t|)\,}$ (28)
Figure 1: The evolution of $H_{num}/H_{p}(\,t-\,D_{2}/p\,)$, where $H_{num}$ is obtained numerically for six different sets of initial conditions. For an attractor solution, the curves must evolve to the value $1$. But, instead, the curves diverge away from $1$, indicating $H$ is not an attractor in time or $H_{p}(\,t-\,D_{2}/p)$ doesn’t exactly represent the evolution of $H$, which is the case here.
@@
where $D_{1}$(a complex value) and $D_{2}$ are integration constants that must be fixed from initial conditions; note that we need two initial values to fix the integration constants completely, in addition to $a_{0}$ and the integration constant $C$, which could be estimated through the time translation symmetry of FRW equations, in 7. From the above equations, note that the condition for an attractor solution is $\lambda^{2}<6$ for an expanding Universe. However, for a contracting Universe, the situation is tricky. The first term inside the square bracket in the equation for $\Delta H$, 27, decays for $\lambda^{2}>6$ in a contracting Universe, and the second term increases; hence, it may appear to diverge. We would like to draw your attention to the fact that this growing term is proportional to $\dot{H}$; that is, this term could be the first term in the Taylor series expansion of H with a time shift $H(t)\approx H_{p}(t)=p/t\rightarrow H_{p}(t-D_{2}/p)=p/(t-D_{2}/p)$, this term, $-D_{2}/p$, owes to the shift in time that appears due to the time translational symmetry of FRW metric. Now, if the analogous term appears in all the physical quantities of the system, like $\phi$, $\dot{\phi}$, et cetera, we can confirm that this term is not physical. To confirm this, one has to calculate the first-order perturbation of $\phi$, and it must contain a term (a growing term for our case), which is $-\dot{\phi}|_{H=H_{p},X=X_{p}}D_{2}/p$. And indeed, this is the case for our scenario. This means the big crunch happens at a different time for different initial conditions. The evolution remains the same other than this time shift. Hence, the growing term is not physical. It was shown for the first time in the famous work [29] and was further extended in [2]. In these papers, the authors calculated the first-order perturbation for the scalar field, $\phi$, and like in our case for the Hubble parameter, they, too, obtained that there is a growing term. They intelligently identified that this term appears due to the time shift, which may be ignored as FRW metrics have time translational symmetry. However, we noticed that our numerical calculation did not match this conclusion. From numerical computation, we noticed that $H$ is not exactly equal to $H_{p}(t-D_{2}/p)$ or evolves to $H_{p}(t-D_{2}/p)$ as required. Instead, it diverges away, see 1, here, $H_{num}/H_{p}(t-D_{2}/p)$ is plotted versus time, $H_{num}$ is the numerically obtained Hubble parameter for different initial conditions. Also, $D_{2}$ is the constant determined from the initial conditions. So, as the initial conditions differ for different curves, both $H_{num}$ and $D_{2}$ are also different. The plot clearly shows that the evolution of $H$ deviates away from $H_{p}(t-D_{2}/p)$ and we must resort to second order calculations to resolve the problem. Let us compute the second-order perturbation of $H$ and $X$.
@@
$+\left(\Delta\dot{X}+\frac{-6X_{p}^{3}-\lambda\sqrt{6}\left(X_{p}^{2}-1\right)+6X_{p}}{2}\Delta\dot{H}\,\kappa\right.$
@@
$+\left(\Delta^{(2)}\dot{X}+\frac{-6X_{p}^{3}-\lambda\sqrt{6}\left(X_{p}^{2}-1\right)+6X_{p}}{2}\Delta^{(2)}\dot{H}\kappa\right.$
$\quad+\frac{-18X_{p}^{2}\Delta X-2\lambda\sqrt{6}X_{p}\Delta X+6\Delta X}{2}\Delta\dot{H}\kappa$
@@
We can separately solve order by order. Solving the second-order equation, we get
$\Delta^{(2)}H=-\frac{{C_{1}}\sqrt{30}}{85t^{\frac{3}{10}}}-\frac{3{D_{1}}{D_{2}}\sqrt{30}}{85t^{\frac{13}{10}}}-\frac{2103t^{\frac{2}{5}}{D_{1}}^{2}}{161840}+\frac{10{D_{2}}^{2}}{t^{3}}+\frac{{C_{2}}}{t^{2}},$ (29)
where the integration constants could be fixed from the initial conditions $\Delta X(t_{i})$ , $\Delta H(t_{i})$, which are the deviations of $H(t_{i})$ and $X(t_{i})$ from $H_{p}(t_{i})$ and $X_{p}(t_{i})$. Now, without loss of generality, we can take $\Delta^{(2)}X(t_{i})=\Delta^{(2)}H(t_{i})=0$; here, $\lambda$ is taken to be $-\sqrt{20}$ since the expression would be cumbersome without assuming a value for $\lambda$.
@@
[…]
$-\frac{3{D_{1}}{D_{2}}\sqrt{30}}{85t^{\frac{13}{10}}}-\frac{2103t^{\frac{2}{5}}{D_{1}}^{2}}{161840}+\frac{{C_{2}}}{t^{2}}+\frac{10\,D_{2}^{2}}{t^{3}}.$ (30)

### v2

@@ Abstract
In this paper, we revisit the stability of power-law models, focusing on an alternative approach that differs significantly from the standard approaches used in studying power-law models. In the standard approach, stability is studied by reducing the system of background FRW equations to a one-dimensional system for a new background variable $X$ in terms of the number of e-foldings. However, we rewrote the equations, incorporating $H$ into the system and went on to do the calculations up to the second order. We demonstrate by computing the deviations from the power-law exact solution to second-order in time and show that power-law contraction is never an attractor in time, regardless of parameter values. Our analysis shows that while first-order corrections align with existing interpretations, second-order corrections introduce significant deviations that cannot be explained by a simple time shift that explains the first-order diverging terms. With importance, we note that in the number of e-folds, the system remains an attractor, while in cosmic time, it is unstable. We also support our claim with numerical results. This new insight has broader implications for the study of attractor behaviour of differential equation solutions and raises questions about the stability of scenarios like the ekpyrotic bounce driven by an exponential potential. Our work also hints that the different temporal variables we use might not be equivalent.
@@ I Introduction
Cosmic inflation and the bounce paradigm represent two pivotal concepts in modern cosmology, offering distinct perspectives on the early universe’s evolution. Proposed by Alan Guth in 1981 [1], cosmic inflation postulates a rapid expansion phase that addresses the horizon problem and the flatness problem. On the other hand, the bounce paradigm challenges the classical Big Bang singularity, proposing a scenario where the universe transitions from a contracting phase to an expanding one without encountering a singularity [2, 3, 4, 5, 6]. This concept has profound implications for cosmological models, providing alternative explanations for the universe’s origins. Bounce cosmology, though less explored compared to inflation, has garnered significant interest among researchers seeking to understand the early universe’s dynamics. Current studies in cosmology emphasize the refinement of both inflationary and bounce models, along with the search for observational evidence to support or refute these paradigms [7, 8]. For a comprehensive understanding of cosmological theories and observations, refer to relevant literature and books [9, 10, 11, 12, 13, 14, 15, 16, 17].
@@
One of the early models of inflation is the power-law inflation proposed by F. Lucchi and S. Matarrese [18]. The model assumes the evolution of the scale factor to be power-law ($a\propto t^{p}$, where $p$ is a positive constant greater than $1$). This is one of the simplest models of inflation that withstood even the latest observational tests [13, 19]. Power-law inflation and its spin-offs [20, 21, 22, 23, 24] are widely studied even today and form the basis of many cosmological models of early Universe, which include models of inflation [25, 22] and cosmic bounce [26, 27], and of late time acceleration [28]. One of the most essential requirements of early Universe models is their ability to solve fine-tuning problems. Hence, a successful model must have the evolution of the Universe to be an attractor. For this reason, the stability analysis of both background evolution and the evolution of cosmological perturbations is important.
In this paper, we reexamine the stability of the background solution of the power-law model. Indeed, it is a topic that has been studied widely in depth and breadth [29, 20, 30, 31, 32, 33, 2, 3, 34]. However, the earlier studies, where they converted the FRW equations in time to equations in number of e-foldings ($N$) for the two quantities defined as $X={\kappa\dot{\phi}}/{(\sqrt{6}H)}$ and $Y=\kappa\sqrt{V}/(\sqrt{3}H)$ [32] missed the point that, in doing so, we might not get the evolution of relevant quantities such as scalar field, Hubble parameter etc., in time. The evolution of these quantities in the number of e-foldings is not enough to talk about the stability of their evolution. Instead, one must look into the evolution of Hubble parameter, scalar field, et cetera in time. Because the number of e-foldings is a function of $H$, the behaviour of $H$ will affect the evolution of the number of e-foldings.
@@
One can note that it is impossible to write these background quantities as a function of the terms $X$ and $Y$. Our analysis, where we consider equations in time, redefines the conditions the parameters must satisfy for a stable expansion and contraction. Previously, there were studies in the literature in this regard, that is, studying the perturbations in time [2, 3]. However, they did the analysis to first-order and came to the conclusion that the perturbation would only result in a shift in time. From our numerical results, we conclude that this is not the case. Further, we analyzed the perturbations to the second order and confirmed our conclusion.
@@
For expansion, that is for an inflationary evolution; though the analysis and results differ, the conclusion regarding the stability of the power-law expansion being an attractor remains the same without any change. However, the conditions for a stable contraction differ drastically. That is, our work is significant for the ekpyrotic bouncing model driven by the exponential potential. Also, our observation that different temporal variables are not completely identical is important.
@@
The outline of this paper is as follows: in the next section, we review the power-law model of inflation and the conditions for its stability. Then, we analytically study the background dynamics, giving importance to the stability of background evolution, where we also compare our numerical results with analytical results. Subsequently, we show that $X$ and other background quantities are attractors in $N$. Finally, in conclusion, we review the results and discuss their importance in the current studies of inflation, bouncing models and stability analysis in general.
@@ II Power-law models
[…]
@@
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa^{2}}R-\frac{1}{2}g^{ab}\nabla_{a}\phi\nabla_{b}\phi-V_{0}e^{\lambda\kappa\left(\phi-\phi_{0}\right)}\right],$ (3)
where general relativity is minimally coupled to a scalar field with an exponential potential with the constants $V_{0}$, $\lambda$ and $\phi_{0}$ as free parameters. The exponential potential allows a smooth inflationary expansion and the cosmological perturbations to have a scale-invariant spectrum. where $t$ is the cosmic time and $a(t)$ is the scale factor. The Friedmann equations are given by
@@
$a(t)=a_{0}(C\,+\,q_{p}\,t)^{p};$ (7a)
$\phi(t)=-\frac{\sqrt{2\,p}\,\ln\,\left(\mathit{q_{p}}t+C\right)}{\kappa}+\phi_{0}$ (7b)
@@
With, $p=2/\lambda^{2}$. We can obtain a contracting solution in two ways: either by assuming $q_{p}$ to be $-1$ and $C$ to take a large positive value, and also here, $t>0$. Now, this can also be achieved by assuming $t$ to take negative values ($\infty$- ∞ to 00), here again, $q_{p}=-1$. In the remaining part of the paper, we take $q_{p}=-1$, $C=0$, and $t$ take negative values for a contracting solution. For an expanding solution $q_{p}=1$, $C=0$ and $t>0$ are assumed.
@@
$\frac{dX}{dN}$ $=-3X\,Y^{2}-\lambda\sqrt{\frac{3}{2}}Y^{2},$ (8)
11 11 $=X^{2}+Y^{2},$ (9)
$\frac{dH}{dN}=-3\,X^{2}\,H,$ (10)
@@
Now, we can see that $X=X_{p}=-\lambda/\sqrt{6}$ and $Y=Y_{p}=\pm\sqrt{1-\lambda^{2}/6}$ is a fixed point, i.e., for this point 8 and 9 are satisfied, and the derivatives of both $X$ and $Y$ are zero. If the evolution exactly falls at these critical points, it remains there indefinitely. Also, these are the very power-law solutions we obtained earlier, 7b, where
@@
$V\left(\phi_{p}\right)$ $=V_{0}\,\left(\mathit{q_{p}}t+C\right)^{-2},$
$\;\;\mbox{where}\;{V_{0}=M_{p}^{2}{q_{p}^{2}}\,p\,\left(3p-1\right)}.$ (15)
@@
We can clearly see that $Y_{p}$ is imaginary for $\lambda^{2}>6$. This is not unexpected as we need a negative value for $V$ in situations like bounce. Also, the value of $Y_{p}$ is not important for our purpose; in the relevant equations and definitions, we don’t have $Y_{p}$; instead, what appears is $Y_{p}^{2}$.
Taking the Taylor expansion of 11 to the first order, we have
@@
So, if $\lambda^{2}<6$, the fixed point solutions are attractor solutions. This argument seems elegant. Also, it was verified numerically for inflationary solutions [29, 20]. For contracting solutions, the ekpyrotic case with $\lambda^{2}>6$ is considered an attractor [6]. For contracting solutions, we can argue that if expanding solutions are repellers, the corresponding contracting solutions would be attractors [41]. That is, $\lambda^{2}>6$ is the required condition for a power-law contraction to be an attractor. In the next section, we revisit the stability analysis.
III Stability of power-law solutions
In the previous section, we obtained fixed-point power-law solutions and the conditions for these solutions to be attractors. However, the quantities that are studied are $X$ and $Y$ and the temporal variable chosen was $N$. But the relevant quantities we have to study in order to see the nature of the evolution of the Universe in this model are $H$, $\phi$ and $\dot{\phi}$ and most importantly, we must study the evolution in time. Also, the analysis gives the conditions for the solutions $X$ and $Y$, defined in the previous section, to be stable when evolved in $N$. Note that those calculations couldn’t explicitly show the evolution of the Hubble parameter and the scalar field in time. Hence, we can argue that the analysis is not complete. That is, we don’t think the stability of the evolution of solutions $X$ and $Y$ in $N$ guarantees that the evolution of $H$, $\phi$ and other relevant quantities will be stable in time. Her, we proceed to reproduce the solutions for perturbed quantities obtained by [2], analytically and then verify them numerically.
@@ III Stability of power-law solutions
$\frac{dX}{dt}=3HX^{3}+\frac{H\sqrt{6}\lambda X^{2}}{2}-\frac{H\sqrt{6}\lambda}{2}-3HX=f_{1}(X,H)$ (21a)
$\frac{dH}{dt}=-3X^{2}H^{2}=f_{2}(X,H)$ (21b)
The relevant solutions, as shown in 16 and 17 are $X=X_{p}$ and the Hubble parameter, which is not a constant in time but a function of time, $H=H_{p}=p/t$ and $X=X_{p}=-\lambda/\sqrt{6}$. Now, Taylor expanding about the exact solution to the first order; we have
$\frac{d}{dt}[X_{p}+\Delta X]$ $=f_{1}(X_{p},H_{p})+\partial_{X}{f_{1}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta{X}+\partial_{H}{f_{1}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta H$ (22)
$\frac{d}{dt}[H_{p}+\Delta H]$ $=f_{2}(X_{p},H_{p})+\partial_{X}{f_{2}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta{X}+\partial_{H}{f_{2}(X,H)|_{\{X=X_{p},H=H_{p}\}}}\Delta H$ (23)
@@
${(\Delta x)}^{.}=\begin{pmatrix}\Delta{x_{1}}\\ \Delta{x_{2}}\end{pmatrix}^{.}\approx\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}\begin{pmatrix}\Delta x_{1}\\ \Delta x_{2}\end{pmatrix}$ (24)
@@
$\Delta x_{1}=\Delta X$, $\Delta x_{2}=\Delta H$ and
@@
$\Delta{X}$ $=$ ${D_{1}}t(-t)^{\frac{-6}{\lambda^{2}}}$ (26)
$\Delta{H}$ $=$ $\left[\frac{2\,D_{1}\sqrt{6}(-t)^{\frac{-6}{\lambda^{2}}}}{\lambda(\lambda^{2}-3)}+\frac{D_{2}}{(-p)}\frac{(-p)}{t^{2}}\right]$ (27)
@@
$\frac{N}{N_{p}}=1+D_{1}\,\frac{\lambda^{3}\sqrt{6}\,t\,(-t)^{\frac{-6}{\lambda^{2}}}}{(\lambda^{4}-9\lambda^{2}+18)\,\mbox{ln}(|t|)}-D_{2}\frac{\lambda^{2}}{2\,t\,\mbox{ln}(|t|)\,}$ (28)
where $N_{p}$ is the expression for the number of e-folding considering $H=H_{p}$. And, $D_{1}$ and $D_{2}$ are integration constants that must be fixed from initial conditions.
@@
Note that in addition to $a_{0}$ and the integration constant $C$ ($a_{0}$ is trivial and $C$ could be estimated through the time translation symmetry of FRW equations, in 7b.) we have two integration constants and need two initial values to fix them completely. From the above equations, note that the condition for an attractor solution is $\lambda^{2}<6$ for an expanding Universe. However, for a contracting Universe, the situation is tricky. The first term inside the square bracket in the equation for $\Delta H$, 27, decays for $\lambda^{2}>6$ in a contracting Universe, and the second term increases; hence, it may appear to diverge. We would like to draw your attention to the fact that this growing term is proportional to $\dot{H}$; that is, this term could be the first term in the Taylor series expansion of H with a time shift $H(t)\approx H_{p}(t)=p/t\rightarrow H_{p}(t-D_{2}/p)=p/(t-D_{2}/p)$, this term, $-D_{2}/p$, owes to the shift in time that appears due to the time translational symmetry of FRW metric. Now, if the analogous term appears in all the physical quantities of the system, like $\phi$, $\dot{\phi}$, $N$, et cetera, we can confirm that this term is not physical. To confirm this, one has to calculate the first-order perturbation of $\phi$, and it must contain a term (a growing term for our case), which is $-\dot{\phi}|_{H=H_{p},X=X_{p}}D_{2}/p$. And indeed, this is the case for our scenario. This means the big crunch happens at a different time for different initial conditions. The evolution remains the same other than this time shift. Hence, the growing term is not physical. It was shown for the first time in [2] and was further extended in [3]. In these papers, the authors calculated the first-order perturbation for the scalar field, $\phi$, and like in our case for the Hubble parameter, they also obtained that there is a growing term. They identified that this term appears due to the time shift, which may be ignored as FRW metrics have time translational symmetry. However, we noticed that our numerical calculation did not match this conclusion. From numerical computation, we noticed that $H$ is not exactly equal to $H_{p}(t-D_{2}/p)$ or evolves to $H_{p}(t-D_{2}/p)$ as required. Instead, it diverges away. Hence, we must resort to second-order calculations to resolve the problem. Let us compute the second-order perturbation of $H$ and $X$.
@@
$+\left(\Delta\dot{X}+\frac{-6X_{p}^{3}-\lambda\sqrt{6}\left(X_{p}^{2}-1\right)+6X_{p}}{2}\Delta H\,\right.$
@@
$+\left(\Delta^{(2)}\dot{X}+\frac{-6X_{p}^{3}-\lambda\sqrt{6}\left(X_{p}^{2}-1\right)+6X_{p}}{2}\Delta^{(2)}H\right.$
$\quad+\frac{-18X_{p}^{2}\Delta X-2\lambda\sqrt{6}X_{p}\Delta X+6\Delta X}{2}\Delta H$
@@
We can separately solve order by order. Solving the second-order equation for $\Delta^{(2)}H$, and assuming $\lambda=-\sqrt{7}$ we get
$\Delta^{(2)}H=$
$\frac{115D_{1}^{2}}{28(-t)^{5/7}}+\frac{3D_{1}D_{2}\sqrt{42}}{14(-t)^{13/7}}-\frac{C_{1}\sqrt{42}}{14(-t)^{6/7}}+\frac{7D_{2}^{2}}{2(t)^{3}}+\frac{C_{2}}{(t)^{2}}$
where the integration constants could be fixed from the initial conditions $\Delta X(t_{i})$, $\Delta H(t_{i})$, which are the deviations of $H(t_{i})$ and $X(t_{i})$ from $H_{p}(t_{i})$ and $X_{p}(t_{i})$ respectively. Now, without loss of generality, we can take $\Delta^{(2)}X(t_{i})=\Delta^{(2)}H(t_{i})=0$; here, $\lambda$ is taken to be $-\sqrt{7}$ since the expression would be cumbersome without assuming a value for $\lambda$. Further, we have assumed $\lambda=-\sqrt{7}$ in the numeric calculations.
@@
$H^{(2)}=\frac{2}{7\,t}$ $+\left(-\frac{\sqrt{7}D_{1}\sqrt{6}}{14(-t)^{6/7}}+\frac{D_{2}}{t^{2}}\right)$
$+\left(\frac{115D_{1}^{2}}{28(-t)^{5/7}}+\frac{3D_{1}D_{2}\sqrt{42}}{14(-t)^{13/7}}\right.$
$\quad\left.-\frac{C_{1}\sqrt{42}}{14(-t)^{6/7}}+\frac{7D_{2}^{2}}{2\,t^{3}}+\frac{C_{2}}{t^{2}}\right)$